OpenStudy (anonymous):

could someone put the work into complete sentences for me? I don't know how id explain it pleasee a_15=3+3(14) a_15=3+42 a_15=45 a_15=2(2)^14 a_15=2(16384) a_15=32768

OpenStudy (anonymous):

What are you trying to explain?

OpenStudy (anonymous):

the work that was done.. just the process but in complete sentences

jimthompson5910 (jim_thompson5910):

it looks like you have 2 formulas: one arithmetic and one geometric

jimthompson5910 (jim_thompson5910):

both are finding the 15th term

OpenStudy (anonymous):

yes! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so let's focus on the arithmetic sequence

jimthompson5910 (jim_thompson5910):

the general formula is an = a1 + d(n-1) where an = nth term a1 = first term d = common difference n = term number

jimthompson5910 (jim_thompson5910):

a_15=3+3(14) is the same as a_15=3+3(15-1), so we can see first term: a1 = 3 common difference: d = 3 term number: n = 15

jimthompson5910 (jim_thompson5910):

we plug those values into an = a1 + d(n-1) and evaluate an = a1 + d(n-1) a15 = 3 + 3(15-1) ... plug in a1 = 3, d = 3, n = 15 a15 = 3 + 3(14) ... subtract a15 = 3 + 42 ... multiply a15 = 45 ... add After you've done those substitutions, you're just using PEMDAS to evaluate

OpenStudy (anonymous):

that's all I have to say for it then?

jimthompson5910 (jim_thompson5910):

there's not much to say really since it's just arithmetic after you plug in those values

jimthompson5910 (jim_thompson5910):

you'd have more to say about the starting term and common difference if anything

OpenStudy (anonymous):

ah! sweet! thanks. so what should I be saying about them?

jimthompson5910 (jim_thompson5910):

I doubt your teacher cares about the arithmetic (in terms of verbal descriptions) just as long as you get the math right you should be fine so you should be talking about the first term and common difference

jimthompson5910 (jim_thompson5910):

basically tell where the sequence starts and how it progresses

jimthompson5910 (jim_thompson5910):

In this case, it starts at 3 and it progresses by adding 3 to each term to get the next one

OpenStudy (anonymous):

okay, well I have the numbers and all the quiz just asks to be written in complete sentences rather than showing the work

jimthompson5910 (jim_thompson5910):

that will lead to the arithmetic nth term formula an = a1 + d(n-1) an = 3 + 3(n-1)

jimthompson5910 (jim_thompson5910):

Ok, then you'll have to describe what's going on instead of writing out the formulas

jimthompson5910 (jim_thompson5910):

so say you're starting with an nth term formula, you plug in the given values and evaluate/compute the result to get 45

jimthompson5910 (jim_thompson5910):

I think there should be a place for the formula though since it kinda speaks for itself

OpenStudy (anonymous):

of course,

jimthompson5910 (jim_thompson5910):

maybe they want full sentences on top of using the formula?

OpenStudy (anonymous):

I guess that could be done. all it tells me is to write the work Id be doing in complete sentences. like I said, I have the formulas, I am just lacking the ability to put what ive done into words

jimthompson5910 (jim_thompson5910):

That's what I'd do honestly if I were in your shoes. I think the formula is important enough that it'd be silly to leave it out. It's probably a case where you have to pretend you're explaining this to a friend or something.

jimthompson5910 (jim_thompson5910):

I guess you could integrate the lines I wrote out into a big paragraph like this maybe We start with the formula an = a1 + d(n-1) and plug in the first term a1 = 3 and common difference d = 3 to get an = 3 + 3(n-1). This is the nth term formula for the given sequence with starting term 3 and common difference 3. To find the 15th term, we plug in n = 15 and evaluate. So a15 = 3 + 3(15-1) = 3+4(14) = 3+42 = 45 is the 15th term of the arithmetic sequence. Something along those lines. That might be what they are aiming for.

OpenStudy (anonymous):

thank you so so much!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

would you be able to help me on the last one that I also solved and has to be written?

jimthompson5910 (jim_thompson5910):

are you able to get started on it?

OpenStudy (anonymous):

on the paragraph?

jimthompson5910 (jim_thompson5910):

yes, what kind of sequence does that formula deal with?

OpenStudy (anonymous):

its pretty much the same thing but with a higher number

jimthompson5910 (jim_thompson5910):

ok what else

OpenStudy (anonymous):

I have the work solved. want to see that?

jimthompson5910 (jim_thompson5910):

I see it at the top (unless you have more work that's not posted?)

OpenStudy (anonymous):

yeah, its a different problem I did

jimthompson5910 (jim_thompson5910):

ok go ahead and post it

OpenStudy (anonymous):

a_20=3+3(19) a_20=3+57 a_20=60 a_20=2(2)^19 a_20=2(524288) a_20=1,048,576

jimthompson5910 (jim_thompson5910):

oh I'm guessing you already wrote the paragraph for a_15=2(2)^14 a_15=2(16384) a_15=32768

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

the first part a_20=3+3(19) a_20=3+57 a_20=60 will follow the same format because this is also an arithmetic sequence (in fact, it has the same starting term 3 and common difference 3, so it's the same sequence)

jimthompson5910 (jim_thompson5910):

the only difference this time is we're finding the 20th term instead of the 15th

jimthompson5910 (jim_thompson5910):

as for a_20=2(2)^19 a_20=2(524288) a_20=1,048,576 this is a geometric sequence with first term of 2 and common ratio of 2 (we're finding the 20th term here too)

OpenStudy (anonymous):

okay, great. I think I know where to go from there then

jimthompson5910 (jim_thompson5910):

alright, I'm glad it's clicking now

OpenStudy (anonymous):

me too! :D I really appreciate your help!

jimthompson5910 (jim_thompson5910):

I'm glad I could help